3.9GHz status

UTC
The Hermitage

The Hermitage

    • 09:00 09:15
      Cavity No. 6 vertical test results 15m
    • 09:15 09:35
      A0 update -- high pressure rinse system, North cave vacuum system, etc. 20m
    • 09:35 09:50
      3rd harmonic helium supply weldments 15m
    • 09:50 10:05
      Repair of cavity No. 2 15m
    • 10:05 10:25
      Other business 20m